# Hallucination to Truth: A Review of Fact-Checking and Factuality Evaluation in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are trained on vast and diverse internet corpora that often include inaccurate or misleading content. Consequently, LLMs can generate misinformation, making robust fact-checking essential. This review systematically analyzes how LLM-generated content is evaluated for factual accuracy by exploring key challenges such as hallucinations, dataset limitations, and the reliability of evaluation metrics. The review emphasizes the need for strong fact-checking frameworks that integrate advanced prompting strategies, domain-specific fine-tuning, and retrieval-augmented generation (RAG) methods. It proposes five research questions that guide the analysis of the recent literature from 2020 to 2025, focusing on evaluation methods and mitigation techniques. The review also discusses the role of instruction tuning, multi-agent reasoning, and external knowledge access via RAG frameworks. Key findings highlight the limitations of current metrics, the value of grounding outputs with validated external evidence, and the importance of domain-specific customization to improve factual consistency. Overall, the review underlines the importance of building LLMs that are not only accurate and explainable but also tailored for domain-specific fact-checking. These insights contribute to the advancement of research toward more trustworthy and context-aware language models.
@@ 1  Introduction
The growing use of Large Language Models (LLMs) in news, healthcare, education, and law means that the accuracy of their content now affects important real-world decisions [1, 2]. These models often generate information that sounds reliable, but can be false or unsupported, increasing the risk of misinformation reaching the public [3]. Because new techniques, datasets, and benchmarks for LLM fact checking are being published at a rapid pace, it has become difficult for researchers and practitioners to keep track of what actually works [4]. This paper provides a systematic review of fact checking in the context of LLMs such as GPT-4 and LLaMA. This review is needed to organize the latest knowledge, highlight the main challenges, and point out areas that require further research. This is especially important at this stage, as LLMs are increasingly shaping the way information is produced, accessed, and trusted throughout society.
Figure 1: The fundamental content structure and categorization of this survey.
@@ 1.1  Challenges
Fact-checking the output of LLM systems faces several intimidating challenges. Among the most notable challenges is the absence of standardized evaluation metrics. Currently used ones quantify surface-level similarity and not factual consistency [5, 6] and are therefore more likely to detect nuanced errors.
Another key limitation of LLMs is hallucination. They tend to produce linguistically consistent but factually inaccurate or entirely fictional text [3]. This is due to language modeling and training on potentially stale or biased data in a data-driven, probabilistic way [7, 8]. Dataset quality is also critical for fact-checking system performance. The majority of benchmarks either have no realistic complexity of real-world claims and are domain independent or are too narrow to be generalized. Furthermore, datasets with unbalanced classes can influence the model response and make the system less robust on a wide spectrum of topics [9].
@@ 1.2  Emerging Innovations
To address these issues, various innovations such as Retrieval-Augmented Generation (RAG), instruction tuning [10, 11], domain-specific fine-tuning [12], multi-agent systems [13, 14], automated self-correction and feedback mechanisms [15, 16], and integration with knowledge graphs [17, 18] have been suggested to overcome these limitations. RAG stands out as a key technique that combines LLMs and external retrieval systems, aligning generated outputs with verifiable sources. RAG architectures have shown notable results in factuality and explainability [19] by allowing LLMs to access and cite external knowledge in real time. These methods are often augmented with advanced prompting techniques, such as hierarchical step-by-step reasoning and multi-agent collaboration [20, 21, 22].
@@ 1.3  Purpose and Research Questions
The objective of the review is to critically evaluate the current prospects of LLM-based fact-checking systems, identify key issues, and investigate the performance of existing solutions. Observing recent growing trends, this paper aims to build more accurate, transparent, and scalable fact-checking systems in LLMs. The review is inspired by five fundamental research questions.
@@
    Rationale: Hallucinations are caused by LLM due to the vast amount of training data containing refined and unverified information. Their output often contains hallucinated answers, which directly affect the trustworthiness and accuracy of the LLM fact check.
@@ 1.4  Contributions
This paper brings three contributions to research on fact-checking using LLMs:
@@
Second. The review combines a wide range of approaches to mitigate hallucinations in LLM output. These range from fine-tuning through domain-specific data to instruction tuning, adversarial training, and self-supervised feedback methods like Self-Checker [20]. In addition, multi-agent architecture and multi-step reasoning strategies are explored as enhancement strategies for factuality and explainability.
Third. The paper offers novel insight into how dataset characteristics such as domain specificity, annotation quality, and multilingual coverage affect the performance of fact checking systems [9].
@@ 1.5  Paper Organization
This paper is organized into eight sections: Section 2 reviews existing research on LLM-based fact checking and highlights key gaps. Section 3 explains the methodology, including how the studies were selected and analyzed. Section 4 presents the findings based on our five core research questions. Section 5 discusses the implications, challenges, and limitations. Section 6 draws attention to open issues and challenges. Section 7 highlights the analysis of future research agendas, and Section 8 concludes the paper with key insights for building more accurate and reliable LLM-based fact-checking systems. Figure 1 illustrates the overall structure of the paper.
Table 1: Comparison of different papers concerning the key RQs. The following RQs drive our review: RQ1: Evaluation Metrics and Gaps; RQ2: Hallucination Effects and Mitigation; RQ3: Datasets and Impact; RQ4: Prompt and Fine-tuning; Domain-specific Training Effects; and RQ5: RAG and Domain-Specific Implementation Challenges.
@@ 2  Related Works
Several researchers have explored fact-checking methods in the context of LLMs. For example, Vykopal et al. [23] conduct a survey of approaches and techniques used in automated fact checking using generative LLMs, such as claim detection, evidence retrieval, and fact verification. They introduce the concept of RAG, which can be used to mitigate challenges such as hallucinations and the use of out-of-date model knowledge utilizing external evidence. However, it does not address the effects of domain-specific training on LLM-based fact checking, the challenges of RAG implementation, or how the quality of the dataset, the specificity of the domain, and the evaluation metrics influence the effectiveness of LLM. Dmonte et al. [24] also explore LLM-based claim verification by analyzing full-system pipelines that include key stages such as evidence retrieval, prompt construction, and explanation generation. They review RAG techniques such as iterative retrieval and claim decomposition, which allows them to address issues such as hallucinations and the challenges of verifying complex or long claims. Evaluation metrics like FactScore and FEVER Score are highlighted to assess and improve factual accuracy. Similarly to [23], this paper overlooks dataset quality, domain-specific challenges, and RAG implementation issues, which are significant in evaluating the reliability of LLM-based fact check.
@@
In another work, Augenstein et al. [3] study the challenges of factual correctness in LLM. They focus on hallucinations, knowledge editing to reduce hallucinations, and the impact of misinformation that AI can spread, which also includes concerns about trust and misuse. They propose some mitigation strategies, such as RAG, although the discussion lacks depth with regard to domain-specific RAG implementations and the associated challenges. Key evaluation metrics for LLM-based fact-checking systems are also discussed to assess factuality, consistency, and text quality, including TruthfulQA, FactScore, GPTScore, G-Eval, SelfCheckGPT, BERTScore, and MoverScore. However, the paper lacks a discussion of several critical technical aspects and challenges, including model interpretability, explainability, and practical implementation and integration of the proposed mitigation techniques in real-world settings. Wang et al. [25] offers a detailed survey of the factuality of LLM, providing a taxonomy of hallucination types and errors in both unimodal and multimodal tasks. A key contribution is mapping factuality challenges to algorithmic solutions and proposing improvements to factuality-aware model calibration. However, the paper does not discuss domain-specific challenges of RAG, prompt design, fact-checking component integration, or system-level architectures and deployment considerations for LLM verification environments.
@@
Although most previous work has briefly discussed evaluation metrics, hallucination effects, and issues related to prompt- or fine-tuning methods, it has not examined the impact of datasets, particularly domain-specific ones. In addition, there is a wide gap between practical challenges and considerations of RAG and domain-specific implementation. Most of the papers gave RAG only a passing reference without discussing larger domain-specific problems. However, our work fully discusses all the primary areas of research questions, such as under-discussed areas, and thus provides a more detailed overview of the field. Table 1 presents a summary of the existing survey papers and shows how they relate to the key research questions of this study.
@@ 3  Methods
To explore how LLMs can be applied to fact-checking, we adopted a structured yet practical approach inspired by well-established research methods [26]. We began by designing a detailed review plan that outlined what we wanted to study, where we would look for relevant literature, and what kind of studies we would include. We then conducted a broad search across leading academic databases, using a combination of manual screening and automated tools to identify the most relevant and up-to-date studies.
The review process consisted of three key phases: (i) planning, (ii) data collection and analysis, and (iii) synthesis and reporting. At the end of our review, we brought together the insights from the selected studies to highlight what is already known, where the gaps are, and what future research should aim to address. By following this clear, step-by-step methodology, we have ensured that our findings are informative and reliable for advancing the use of LLMs in fact-checking tasks [27].
@@ 3.1  Search Strategy
To ensure comprehensive coverage of relevant literature, we designed a focused search strategy using well-defined keywords and Boolean operators. Our queries were tailored to capture publications related to the use of LLMs in fact-checking and related tasks. The following keyword combinations were used across databases such as IEEE Xplore, ACM Digital Library, Wiley Online Library, ScienceDirect, Web of Science, arXiv, and Google Scholar:
  • "large language models" AND "fact-checking"
  • "LLM" AND "misinformation detection"
  • "automated fact verification" AND "LLMs"
  • "factuality evaluation" AND "natural language processing"
  • "hallucination" AND "LLMs"
  • "LLM hallucination"
  • "hallucination mitigation" AND "large language models"
  • "hallucination detection" AND "large language models"
  • "fact-checking datasets" OR "benchmark datasets for fact verification"
  • "retrieval-augmented generation" AND "fact-checking"
  • "rag" AND "LLM"
  • "rag" AND "fact-checking"
  • "fine-tuning" AND "fact verification models"
  • "prompt engineering" AND ("truthful generation" OR "fact-checking")
  • "LLM-based fact verification" AND "NLP"
Our selection process targeted significant areas, including fact-checking methods, model assessment techniques, dataset analysis considerations, and optimization procedures. Table 2 reports the details of the libraries, along with the number of publications we selected for further analysis.
Table 2: Number of articles selected from online libraries for further analysis.
No. Library Quantity
1 IEEE Xplore 635
2 ACM Digital Library 769
3 Scopus 809
4 Wiley Online Library 153
5 Web of Science 281
6 Others 997
Total 3644
@@ 3.2  Selection Criteria
Each article is carefully assessed using our evaluation criteria to determine whether it meets the inclusion or exclusion requirements. The key inclusion criteria (IC) and exclusion criteria (EC) are shown in Figure 2.
Figure 2: Inclusion and exclusion criteria for article selection.
@@ 3.3  Article Selection
We conducted a systematic review of articles from leading conferences and journals at the intersection of fact-checking, Natural Language Processing (NLP), and LLMs. An overview of the number of selected journals, conference proceedings, and preprints is shown in Figure 3.
@@
The review focused on studies published between 2020 and 2025 that employed LLMs to verify external claims or factual content, excluding works solely analyzing hallucinations or internal factual consistency. From an initial pool of 3,644 records retrieved from various academic databases, we applied a multi-stage screening process, comprising duplicate removal, title and abstract screening, and full-text evaluation. Finally, we have selected 57 articles that meet our inclusion criteria and align with the objectives of this review. The article selection workflow is illustrated in Figure 4, which outlines the stages of identification, screening, eligibility, and final inclusion.
Figure 4: The article selection and screening process of this survey.
@@ 4  Findings from the Research Questions
In this section, we present a comprehensive key result finding focusing on evaluation metrics, the impact of hallucinations in LLMs, datasets, prompt designing and fine-tuning, and integration of RAG.

### v2

@@ Abstract
Large Language Models (LLMs) are trained on vast and diverse internet corpora that often include inaccurate or misleading content. Consequently, LLMs can generate misinformation, making robust fact-checking essential. This review systematically analyzes how LLM-generated content is evaluated for factual accuracy by exploring key challenges such as hallucinations, dataset limitations, and the reliability of evaluation metrics. The review emphasizes the need for strong fact-checking frameworks that integrate advanced prompting strategies, domain-specific fine-tuning, and retrieval-augmented generation (RAG) methods. It proposes five research questions that guide the analysis of the recent literature from 2020 to 2025, focusing on evaluation methods and mitigation techniques. Instruction tuning, multi-agent reasoning, and RAG frameworks for external knowledge access are also reviewed. The key findings demonstrate the limitations of current metrics, the importance of validated external evidence, and the improvement of factual consistency through domain-specific customization. The review underscores the importance of building more accurate, understandable, and context-aware fact-checking. These insights contribute to the advancement of research toward more trustworthy models.
@@ 1  Introduction
The growing use of Large Language Models (LLMs) in news, healthcare, education, and law means that their accuracy directly affects real-world decisions [1, 2]. These models often generate reliable information but may be false, which poses a risk of misinformation [3]. As newer fact-checking techniques, datasets, and benchmarks are published rapidly, it is challenging for researchers and practitioners to keep track of effective solutions [4]. This paper presents a systematic review of fact-checking related to LLMs, organizing recent advancements in this field, identifying and categorizing ongoing challenges, and highlighting potential avenues for future research. This is crucial as LLMs increasingly shape the way information is trusted and used in society.
Figure 1: The fundamental content structure and categorization of this survey. The framework outlines eight sections, beginning with introduction, related works, and methods, followed by results addressing five research questions on evaluation metrics, hallucinations, datasets, prompting strategies, and retrieval-augmented generation. Subsequent sections cover discussion, open challenges, future research agendas, and conclusion, providing a coherent roadmap of the survey.
@@ 1.1  Challenges
Fact-checking the output of LLM systems faces several intimidating challenges. The lack of standardized evaluation metrics is one of the most notable challenges. Currently used ones quantify surface-level similarity rather than factual consistency [5, 6] and are therefore less effective at detecting nuanced errors.
Another key limitation of LLMs is hallucination. They tend to produce linguistically consistent but factually inaccurate or entirely fictional text [3]. This occurs due to language modeling and training on potentially stale or biased data using the probabilistic method [7, 8]. Dataset quality is also critical for fact-checking system performance. The majority of benchmarks either lack the realistic complexity of real-world claims and are domain-independent or are too narrow to be generalized. Furthermore, datasets with imbalanced classes can influence the model response and make the system less robust on a wide spectrum of topics [9].
@@ 1.2  Emerging Innovations
To address these issues, various innovations have been suggested, including Retrieval-Augmented Generation (RAG), instruction tuning [10, 11], domain-specific fine-tuning [12], multi-agent systems [13, 14], automated self-correction and feedback mechanisms [15, 16], and integration with knowledge graphs [17, 18]. RAG stands out as a key technique that combines LLMs and external retrieval systems, aligning generated outputs with verifiable sources. RAG architectures have shown notable results in factuality and explainability [19] by allowing LLMs to access and cite external knowledge in real time. These methods are often augmented with advanced prompting techniques, such as hierarchical step-by-step reasoning and multi-agent collaboration [20, 21, 22].
@@ 1.3  Purpose and Research Questions
The objective of the review is to critically evaluate the current prospects of LLM-based fact-checking systems, identify key issues, and investigate the performance of existing solutions. Observing recent growing trends, this paper aims to build more accurate, transparent, and scalable fact-checking systems in LLMs. The review is inspired by five fundamental research questions (RQ).
@@
    Rationale: Hallucinations are caused by LLM due to the vast amount of training data containing refined and unverified information. Their output often contains hallucinated answers, which directly affect the trustworthiness and accuracy of the LLM fact-checking.
@@ 1.4  Contributions
This paper presents three contributions to research on fact-checking using LLMs:
@@
Second. The review combines a wide range of approaches to mitigate hallucinations in LLM output. These range from fine-tuning through domain-specific data to instruction tuning, adversarial training, and self-supervised feedback methods like Self-Checker [20]. In addition, multi-agent architecture and multi-step reasoning techniques are explored to enhance factuality and explainability.
Third. The paper offers novel insight into how dataset characteristics such as domain specificity, annotation quality, and multilingual coverage affect the performance of fact-checking systems [9].
@@ 1.5  Paper Organization
This paper is organized into eight sections: Section 2 reviews existing research on LLM-based fact-checking and highlights key gaps. Section 3 explains the methodology, including how the studies were selected and analyzed. Section 4 presents the findings based on our five core research questions. Section 5 discusses the implications, challenges, and limitations. Section 6 draws attention to open issues and challenges. Section 7 highlights the analysis of future research agendas, and Section 8 concludes the paper with key insights for building more accurate and reliable LLM-based fact-checking systems. Figure 1 illustrates the overall structure of the paper.
Table 1: Comparison of different papers concerning the key RQs. The following RQs guide our review: RQ1: Evaluation Metrics and Gaps; RQ2: Hallucination Effects and Mitigation; RQ3: Datasets and Impact; RQ4: Prompt and Fine-tuning; Domain-specific Training Effects; and RQ5: RAG and Domain-Specific Implementation Challenges.
@@ 2  Related Works
Several researchers have explored fact-checking methods in the context of LLMs. For example, Vykopal et al. [23] conduct a survey of approaches and techniques used in automated fact-checking using generative LLMs, such as claim detection, evidence retrieval, and fact verification. They introduce the concept of RAG, which can be used to mitigate challenges such as hallucinations and the use of out-of-date model knowledge, using external evidence. However, it does not address the effects of domain-specific training on LLM-based fact-checking, the challenges of RAG implementation, or how the quality of the dataset, the specificity of the domain, and the evaluation metrics influence the effectiveness of LLMs. Dmonte et al. [24] also explore LLM-based claim verification by analyzing full-system pipelines that include key stages such as evidence retrieval, prompt construction, and explanation generation. They review RAG techniques, including iterative retrieval and claim decomposition, which allow them to address issues such as hallucinations and the challenges of verifying complex or long claims. Evaluation metrics like FactScore and FEVER Score are highlighted to assess and improve factual accuracy. Similar to [23], this paper overlooks dataset quality, domain-specific challenges, and RAG implementation issues, which are significant in evaluating the reliability of LLM-based fact-checking.
@@
In another work, Augenstein et al. [3] study the challenges of factual correctness in LLM. They focus on hallucinations, knowledge editing to reduce hallucinations, and the impact of misinformation that AI can spread, which also includes concerns about trust and misuse. They propose some mitigation strategies, such as RAG, although the discussion lacks depth on domain-specific RAG implementations and the associated challenges. Key evaluation metrics for LLM-based fact-checking systems are also discussed to assess factuality, consistency, and text quality, including TruthfulQA, FactScore, GPTScore, G-Eval, SelfCheckGPT, BERTScore, and MoverScore. However, the paper lacks a discussion of several critical technical aspects and challenges, including model interpretability, explainability, and practical implementation and integration of the proposed mitigation techniques in real-world settings. Wang et al. [25] offer a detailed survey of the factuality of LLM, providing a taxonomy of hallucination types and errors in both unimodal and multimodal tasks. A key contribution is mapping factuality challenges to algorithmic solutions and proposing improvements to factuality-aware model calibration. However, the paper does not discuss domain-specific challenges of RAG, prompt design, fact-checking component integration, or system-level architectures and deployment strategies for LLM verification environments.
@@
Although most previous work has briefly discussed evaluation metrics, hallucination effects, and issues related to prompt- or fine-tuning methods, it has not examined the impact of datasets, particularly domain-specific ones. In addition, there is a wide gap between practical challenges and considerations of RAG and domain-specific implementation. Most of the papers give RAG only a passing reference without discussing larger domain-specific problems. However, our work thoroughly addresses the primary research questions, especially the under-discussed areas, and thus provides a more detailed overview of the field. Table 1 presents a summary of the existing survey papers and shows how they relate to the key research questions of this study.
@@ 3  Methods
To explore how LLMs can be applied to fact-checking, we adopted a structured and practical approach inspired by well-established research methods [26]. The review process consisted of three key phases: (i) planning, (ii) data collection and analysis, and (iii) synthesis and reporting. Firstly, we defined the overall scope of the review and ensured alignment with our RQs (i.e., RQ1-RQ5). This included developing a detailed review protocol that specified the objectives, inclusion and exclusion criteria, and the databases to be searched. We also identified the dimensions most relevant to LLM fact-checking, including evaluation metrics, hallucinations, datasets, prompting and fine-tuning methods, and RAG, and set these as the guiding categories for further analysis.
We then performed a broad search across leading academic databases, using a combination of manual screening and automated tools to identify the most relevant and up-to-date studies. Following this, we have analyzed the studies according to the methodological approach employed (i.e., benchmarking, prompting strategy, dataset evaluation), which RQ it addressed, and the contribution type (e.g., framework, metric, dataset, or application). At the end of our review, we brought together these insights from the selected studies to highlight what is already known, where the gaps are, and what future research should aim to address. By following this clear and step-by-step methodology, we have ensured that our findings are informative and reliable to advance the use of LLM in fact-checking tasks [27]. The following sections detail the process in each aspect.
@@ 3.1  Search Strategy
To ensure a comprehensive review of the relevant literature, we developed a focused search strategy employing well-defined keywords and Boolean operators. Our objective was to capture publications related to the application of LLMs in fact-checking and associated tasks. The search was conducted across several major academic databases, including Web of Science, arXiv, Scopus, and OpenReview, and covered a publication window from January 1, 2021, to September 15, 2025. For the Scopus and Web of Science databases, we utilized the following comprehensive search query:
("large language models" AND "fact-checking") OR ("LLM" AND "misinformation detection") OR ("automated fact verification" AND "LLMs") OR ("factuality evaluation" AND "natural language processing") OR ("hallucination" AND "LLMs") OR ("LLM hallucination") OR ("hallucination mitigation" AND "large language models") OR ("hallucination detection" AND "large language models") OR ("fact-checking datasets" OR "benchmark datasets for fact verification") OR ("retrieval-augmented generation" AND "fact-checking") OR ("RAG" AND "LLM") OR ("RAG" AND "fact-checking") OR ("fine-tuning" AND "fact verification models") OR ("prompt engineering" AND ("truthful generation" OR "fact-checking")) OR ("LLM-based fact verification" AND "NLP") OR (("misinformation detection" OR "fake news") AND "hallucination")
The specific search keys used for OpenReview and arXiv are detailed in our publicly available GitHub repository. Our selection process was designed to target significant research areas, including fact-checking methodologies, model assessment techniques, dataset analysis, and optimization procedures. The initial search produced the results summarized in Table 2.
Table 2: Number of Publications Retrieved from Each Database
Database Total Papers Retrieved
Web of Science 1,235
arXiv 9,180
Scopus 3,270
OpenReview ~1,000
Total 14,685
The high number of results from arXiv is primarily a consequence of its search limitations. The platform does not handle complex, multi-term queries well, which requires us to search for each keyword separately. This method inevitably led to the same articles being counted multiple times, explaining the inflated total.
@@ 3.2  Selection Criteria
Each article is carefully assessed using our evaluation criteria to determine whether it meets the inclusion or exclusion requirements. The key inclusion criteria (IC) and the exclusion criteria (EC) are shown in Figure 2.
Figure 2: The article selection and screening process is on the left, and inclusion and exclusion criteria for article selection are on the right.
@@ 3.3  Article Selection
We conducted a systematic review of articles from leading conferences and journals at the intersection of fact-checking, Natural Language Processing (NLP), and LLMs. The review focused on studies published between January 1, 2021, and September 15, 2025, that employed LLMs to verify external claims or factual content, excluding works that solely analyzed hallucinations or internal factual consistency.
From an initial pool of 14,685 records retrieved from various academic databases, we applied a multi-stage screening process, comprising duplicate removal, title and abstract screening, and full-text evaluation. The article selection workflow is illustrated in Figure 2, which outlines the stages of identification, screening, eligibility, and final inclusion. Finally, we selected 64 articles that meet our inclusion criteria and align with the objectives of this review, with publication dates ranging from September 2022 to August 2025.
An overview of the number of selected journals, conference proceedings, and preprints is shown in Figure 3. Figure 4 demonstrates the monthly publication frequency of the selected articles, with the majority published in 2024 and 2025. Furthermore, Figure 5 summarizes the geographical distribution of the publications, indicating that regions with well-established research ecosystems and institutional support contributed the largest share of works, which often correlates with higher representation in prestigious venues and greater citation visibility.
@@
Figure 4: Grouped bar chart showing the monthly breakdown of publications from 2021 to 2025.
Figure 5: Bar chart illustrating the geographical breakdown of publications.
To ensure a comprehensive scope, we examined the distribution of publication venues. Due to the fact that various channels influence methodological priorities, venue analysis is crucial. RAG, prompting techniques, and assessment measures are commonly highlighted at prestigious NLP conferences, including ACL, EMNLP, and NeurIPS. Understanding publication frequency in conjunction with venue dispersion sheds light on the field’s structural and temporal dynamics. Because it has a direct impact on methodological variety and research impact, venue concentration is important. Peer review procedures at high-impact conferences and publications are usually more stringent, guaranteeing methodological originality, more robust empirical validation, and wider recognition. Because of this, work that is published in these types of forums typically receives more attention and citations, which supports the prevailing paradigms in the area. This pattern also reflects how research environments with greater institutional support and global visibility tend to secure stronger representation in prestigious venues, which in turn shapes methodological diversity and amplifies citation impact. Domain-specific journals, on the other hand, provide important diversity by publishing contextually grounded research that may not meet the innovation-focused requirements of prestigious venues, despite the fact that they are frequently less referenced.
@@ 4  Findings from the Research Questions
In this section, we present a comprehensive key result finding focusing on evaluation metrics, the impact of hallucinations on LLMs, datasets, prompt design and fine-tuning, and integration of RAG.